# Enhancing Fatigue Detection through Heterogeneous Multi-Source Data Integration and Cross-Domain Modality Imputation

### v3

@@ Abstract
Fatigue detection for human operators plays a critical role in safety critical applications such as aviation, mining, and long haul transport. While numerous studies have demonstrated the effectiveness of high-fidelity sensors in controlled laboratory environments, their performance often degrades when ported to real-world settings due to noise, lighting condition, and field of view constraints, thereby limiting their practicality. This paper formalizes a deployment-oriented setting for real-world fatigue detection, where high-quality sensors are often unavailable in practical applications. To address this challenge, we propose leveraging knowledge from heterogeneous source domains, including high-fidelity sensors that are difficult to deploy in the field but commonly used in controlled environments, to assist fatigue detection in the real-world target domain. Building on this idea, we design a heterogeneous and multi-source fatigue-detection framework that adaptively utilizes the available modalities in the target domain while exploiting diverse configurations in the source domains through cross-domain alignment and modality imputation. Our experiments, conducted using a field-deployed sensor setup and two publicly available human fatigue datasets, demonstrate the practicality, robustness, and improved generalization of our approach across subjects and domains. The proposed method achieves consistent gains over strong baselines in sensor constrained scenarios.
Note to Practitioners—The proposed work is motivated by the growing need for practical and reliable fatigue detection in industrial applications where operator fatigue poses serious safety and productivity risks. While high-performing physiological sensors like EEG provide high-fidelity signals, their use in these environments is limited by constraints such as cost, setup complexity, and sensitivity to environmental conditions. Thus, industry deployment often relies on context-appropriate sensors that may not fully capture fatigue related states. To resolve this gap, we propose a multi-source heterogeneous fatigue detection framework that enables field-deployed systems to take advantage of high-fidelity data collected from controlled environments, even if those modalities are unavailable at deployment. This is achieved through cross-domain alignment and modality imputation. Our experiments validate the practicality, robustness, and generalization of the proposed approach, showing consistent improvements in subject-independent and cross-domain settings.
@@ I Introduction
Although the aforementioned approaches address the challenge of sensor unreliability through data augmentation, few efforts have tackled the equally critical issue of sensor accessibility. In many real-world scenarios, high-end sensors, such as EEG, imaging systems, and pupil dilation monitors, while reliable and accurate, remain prohibitively expensive. The extensive calibration, routine maintenance, and specialized operator training required make these sensors difficult to operate for wide, real-world deployment. Moreover, these sensors are highly sensitive to environmental factors, like noise, lighting condition, and field of view constraints. These sensitivities complicate data collection and analysis in real-world scenarios, ultimately degrading model accuracy and robustness [e19080385]. As a result, studies involving these sensors are often confined to simulated and controlled environments. For example, high resolution EEG was deployed in a simulated environment, together with heart rate and eye blink rate, to analyze mental fatigue or drowsiness during car driving [6347469]. However, such simulation based setups often fail to capture the complexity and uncertainty of actual driving conditions [8482470] , which in turn limits the generalizability of models trained in these idealized environments. Nonetheless, transferring knowledge gained from controlled settings remains valuable, provided it is adapted appropriately, offering a foundation for practical fatigue detection in real-world scenarios where direct sensor deployment is constrained. Note that this type of transfer requires a form of cross-domain adaptation that goes beyond the capabilities of existing generative methods. Current approaches, such as GANs and GPT-based models mentioined earlier, have proven effective for within-domain synthesis, where additional data is generated for modalities already present at both training and deployment (e.g., generating more EEG from partial EEG). However, they are not designed for cross-domain transfer with target-never-seen modalities, where an entire sensor modality is absent in the target domain at both training and deployment, for example, leveraging EEG-related knowledge when no EEG has ever been observed in the target system. This limitation highlights a critical gap: the need for methods that can support cross-domain sensor transfer to address real-world sensor accessibility constraints.
@@
Motivated by these remarks, this paper proposes a novel domain-adaptive framework that enables real-world fatigue detection systems to learn from and be enhanced by controlled-environment data, even when the data originate from entirely different sensor sources. This approach offers a practical solution by facilitating cross-domain knowledge transfer, extending the utility of high-fidelity but deployment-constrained sensors to settings where they are otherwise unavailable. To this end, we present the framework with the following core contributions:
@@ II Related Work
This section reviews key developments in fatigue assessment. Section II.A examines existing fatigue detection methods, highlighting their advantages and limitations, which underscore the need for a new approach that builds on strengths while addressing shortcomings. Section II.B explores the current data generation methods in mitigating data scarcity, along with their constraints in practical applications. Finally, Section II.C discusses the use of open datasets for fatigue assessment and the associated challenges with sensor discrepancies and label heterogeneity.
@@ II-A Fatigue Assessment
Traditional fatigue assessment methods are based on subjective evaluation tools, including self-reports and questionnaires. Since the last century, numerous classic fatigue assessment scales have been widely utilized, including the Stanford Sleepiness Scale (SSS) [Shahid2012SSS], proposed by Stanford University; the Multidimensional Fatigue Inventory (MFI) [smets1995multidimensional], developed by Smets et al.; and the International Fitness Scale (IFIS) [Ortega2011], which measures changes in overall vigor and affect. These scales encompass a multitude dimensions of fatigue and serve as the foundation for traditional fatigue assessment. Nevertheless, both questionnaires and self-reports are highly influenced by an individual’s psychological state and external environment, resulting in subjective and inconsistent outcomes [gawron2016overview]. Moreover, most of these methods [brown2016fatigue, jap2009using, kar2010eeg] are post hoc assessments, making them unsuitable for real-time fatigue monitoring and preventive measures-both of which are essential in high-stakes environments, such as driving [app14073016] and aviation [article111].
@@
Among various physiological signals, EEG has gained popularity in fatigue detection due to its ability to directly measure brain activity with high accuracy. However, traditional EEG systems are expensive, require an intrusive setup and specialized personnel for operation, making it challenging for real-world deployment. As a result, the majority of EEG-based fatigue studies remain confined to controlled laboratory settings [trejo2015eeg, cheng2011mental, monteiro2019using]. Although portable EEG devices have emerged for field applications, their data quality is not as good as that of laboratory-grade devices. Additionally, they are susceptible to external noise and interference, significantly impacting their stability and detection accuracy [larocco2020systemic, shashidhar2020real].
@@ II-B Data Synthesis Methods
Deep learning, on the other hand, has provided promising solutions to this challenge. Recurrent Neural Networks (RNNs), particularly Long Short-Term Memory (LSTM) networks, have proven effective in capturing long-term temporal dependencies in multivariate time-series data. Several studies have demonstrated their success in a variety of applications, such as GRU-D, which incorporates decay mechanisms and masking for clinical time-series data [che2016recurrent]; an LSTM-based imputation model for ECG signals that outperforms classical methods [verma2019accurate]; and BRITS, which enables bidirectional gradient-based imputation without strong distributional assumptions [cao2018brits]. In parallel, Convolutional Neural Networks (CNN) have shown great success in extracting local spatial and temporal features. Their strength has inspired hybrid models that combine CNNs with LSTMs to jointly leverage spatial and temporal dependencies for more accurate imputation and signal reconstruction [Eum2022Imputation]. Additionally, GANs have emerged as a leading approach for producing high-fidelity synthetic signals that are statistically indistinguishable from real ones. Examples can be found in [10107629, 10239214][NIA2024110129].
@@ II-C Open Fatigue Detection Datasets
In addition to these sensor-level discrepancies, label heterogeneity also presents a major challenge for joint modeling and cross-dataset transfer. For instance, some datasets use continuous fatigue scores, whereas others rely on binary labels such as “fatigued” or “not fatigued.” Even among continuous labels, threshold definitions may differ (e.g., fatigue defined as scores >3>3 or >5>5 on a subjective scale). These inconsistencies need to be addressed to enable more consistent model transfer across datasets.
@@ III-D Framework
$\mathcal{X}^{s-}=\mathcal{X}_{\text{T}}\setminus\mathcal{X}_{s},\qquad\mathcal{X}^{s+}=\mathcal{X}_{s}\setminus\mathcal{X}_{\text{T}}$
@@
3:  Apply $g_{s}$ to impute missing features $\mathcal{X}^{s+}$ in $\mathcal{D}_{T}$ based on $\mathcal{D}_{s}$
@@ IV-A Experimental Setup
For the experiments to be elaborated in a later subsection, three datasets containing various physiological signals some overlapping and others unique are selected. Meanwhile, three widely used classification methods are chosen, utilizing both individual datasets and the enhanced target data for training. To minimize discrepancies and enhance the usability of these datasets, data pre-processing is required before prior to experimentation. To that end, we describe the selected datasets, implementation environment, and the data pre-processing steps in detail.
@@ IV-A2 Implementation Environment and Data Preprocessing
For all experiments described below, except for the sensitivity test, the learning parameters are initialized as listed in Table II.
@@ IV-B3 Impact of Cross-Domain Imputed Modalities on Classification Performance
To assess the robustness and generalizability of our approach across model architectures, we train four different classifiers in this experiment—MLP, LSTM, CNN1D, and a Transformer—on each dataset variant. We assess performance using multiple metrics, including accuracy, binary F1, Area Under the ROC Curv(ROC AUC), Area Under the Precision–Recall Curve(PR AUC), macro F1, weighted F1, expected calibration error (ECE), and Brier score. This setup allows us to evaluate not only how each additional reconstructed modality—individually or in combination—contributes to downstream fatigue detection performance, but also the effectiveness of our framework across diverse classifier types.
@@ IV-B4 Ablation of Batch Normalization and Jacobian Regularization
((a))
((b))
Figure 4: Sensitivity analysis of the Transformer on VPFD (EEG+ECG).
IV-B5 Sensitivity Tests
Building on the previous results, where the combination of TimeMixer as the imputer and Transformer as the classifier achieved the better gains, we perform sensitivity tests under this configuration. Specifically, we evaluate the Transformer on VPFD (EEG + ECG) augmented using both FatigueSet and MEFAR via TimeMixer. The goal is to assess model robustness to input perturbations and hyperparameter variations (Fig.4).
For noise robustness, we inject zero-mean Gaussian noise with varying intensity $\sigma$ into the standardized input features and fit a low-order regression of validation AUC against $\sigma$. The resulting curve remains nearly flat for small $\sigma$ values and decreases smoothly as $\sigma$ increases, indicating stability under mild perturbations and controlled degradation under stronger noise (Fig.4a).
For regularization, we vary the affine parameters $(\alpha,\beta)$ applied to standardized inputs as $x^{\prime}=\alpha x+\beta$, and fit a smooth response surface of AUC over $(\alpha,\beta)$. The surface exhibits a clear ridge at mid-range values, with performance declining only when $\alpha$ or $\beta$ deviates substantially from their nominal values (e.g., $\alpha>1$ or $|\beta|>0.2$), indicating robustness to moderate calibration differences but sensitivity to extreme scaling or shifting. The results indicate that the model remains stable under small scaling and bias variations, which are typical in real-world deployments where sensor calibration and preprocessing may differ.
@@ V Conclusion
Fatigue detection is crucial in high stakes environments, where reduced alertness can have serious consequences. Despite advances in sensor technology, reliability and accessibility in real world environments remain major challenges, as many high-end sensors, such as EEG, are expensive, highly sensitive to environmental conditions, and require substantial operational setup outside controlled settings. To address this challenge, this paper proposes a proof-of-concept framework that enables knowledge transfer across differently instrumented sensor domains, aiming to enhance fatigue detection in real-world sensor-constrained scenarios. In particular, our approach transfers knowledge learned from high-fidelity laboratory sensors to field-deployable systems by using shared-channel overlap as a bridge, thereby maintaining detection accuracy when some modalities are unavailable. Our experiments provide empirical support for our claim and validate the framework’s effectiveness and practicality, showing that transferred knowledge can substantially enhance detection when high-end sensors are unavailable.
The research can be extended in several directions. Although label heterogeneity is not present in our current case study, it remains an important factor for generalization in real-world applications. Differences in annotation subjectivity can introduce transfer bias and evaluation challenges. Thus approaches incorporating open-set recognition and label-shift calibration [10.1145/2768831] are worth exploring to mitigate such bias. In addition, our experiments show that using TimerMixer and Transformer yields small but consistent gains over other traditional configurations. We will further this exploration in our future research, developing jointly optimized imputation and classifier architectures to better capture nonlinear dependencies and enhance fatigue-detection performance. Finally, as the framework continues to evolve, we plan to examine its real-time deployment aspects, focusing on computational efficiency, latency, and resource requirements to ensure suitability for field applications.

### v4

@@ Abstract
Fatigue detection for human operators plays a critical role in safety critical applications such as aviation, mining, and long haul transport. While numerous studies have demonstrated the effectiveness of high-fidelity sensors in controlled laboratory environments, their performance often degrades when ported to real-world settings due to noise, lighting condition, and field of view constraints, thereby limiting their practicality. This paper formalizes a deployment-oriented setting for real-world fatigue detection, where high-quality sensors are often unavailable in practical applications. To address this challenge, we propose leveraging knowledge from heterogeneous source domains, including high-fidelity sensors that are difficult to deploy in the field but commonly used in controlled environments, to assist fatigue detection in the real-world target domain. Building on this idea, we design a heterogeneous and multi-source fatigue-detection framework that adaptively utilizes the available modalities in the target domain while exploiting diverse configurations in the source domains through cross-domain alignment and modality imputation. As a proof-of-concept, we intentionally employ well-established algorithms for both knowledge transfer and fatigue detection components to validate the framework’s core idea. Experiments using a field-deployed sensor setup and two publicly available human fatigue datasets, demonstrate the practicality, robustness, and improved generalization of our approach, achieving consistent gains over strong baselines in sensor constrained scenarios.
@@
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Although the aforementioned approaches address the challenge of sensor unreliability through data augmentation, few efforts have tackled the equally critical issue of sensor accessibility. In many real-world scenarios, high-end sensors, such as EEG, imaging systems, and pupil dilation monitors, while reliable and accurate, remain prohibitively expensive. The extensive calibration, routine maintenance, and specialized operator training required make these sensors difficult to operate for wide, real-world deployment. Moreover, these sensors are highly sensitive to environmental factors, like noise, lighting condition, and field of view constraints. These sensitivities complicate data collection and analysis in real-world scenarios, ultimately degrading model accuracy and robustness [e19080385]. As a result, studies involving these sensors are often confined to simulated and controlled environments. For example, high resolution EEG was deployed in a simulated environment, together with heart rate and eye blink rate, to analyze mental fatigue or drowsiness during car driving [6347469]. However, such simulation-based setups often fail to capture the complexity and uncertainty of actual driving conditions [8482470], which in turn limits the generalizability of models trained in these idealized environments. Nonetheless, transferring knowledge gained from controlled settings remains valuable, provided it is adapted appropriately, offering a foundation for practical fatigue detection in real-world scenarios where direct sensor deployment is constrained. Note that this type of transfer requires a form of cross-domain adaptation that goes beyond the capabilities of existing generative methods. Current approaches, such as GANs and GPT-based models mentioined earlier, have proven effective for within-domain synthesis, where additional data is generated for modalities already present at both training and deployment (e.g., generating more EEG from partial EEG). However, they are not designed for cross-domain transfer with target-never-seen modalities, where an entire sensor modality is absent in the target domain at both training and deployment, for example, leveraging EEG-related knowledge when no EEG has ever been observed in the target system. This limitation highlights a critical gap: the need for methods that can support cross-domain sensor transfer to address real-world sensor accessibility constraints.
@@
Motivated by these remarks, this paper proposes a novel domain-adaptive framework that enables real-world fatigue detection systems to learn from and be enhanced by controlled-environment data, even when the data originate from entirely different sensor sources. By facilitating cross-domain knowledge transfer, the proposed approach extends the utility of high-fidelity but deployment-constrained sensors to practical settings where they are otherwise unavailable. Importantly, this work serves as a proof-of-concept, aiming to demonstrate the feasibility and effectiveness of the proposed heterogeneous and multi-source fatigue-detection framework. To isolate the contribution of the framework design itself, we intentionally employ widely used, well-understood models for both the knowledge transfer and fatigue detection components, emphasizing that the innovation lies in their integration rather than in the specific algorithms themselves. The proposed framework is presented with the following key contributions:
@@ II Related Work
This section reviews key developments in fatigue assessment. Section II.A examines existing fatigue detection methods, highlighting their advantages and limitations, which underscore the need for a new approach that builds on strengths while addressing shortcomings. Section II.B explores the current data generation methods in mitigating data scarcity, along with their constraints in practical applications. Finally, Section II.C discusses the use of open datasets for fatigue assessment and the associated challenges with sensor discrepancies.
@@ II-A Fatigue Assessment
Traditional fatigue assessment methods are based on subjective evaluation tools, including self-reports and questionnaires. Since the last century, numerous classic fatigue assessment scales have been widely utilized, including the Stanford Sleepiness Scale (SSS) [Shahid2012SSS], proposed by Stanford University; the Multidimensional Fatigue Inventory (MFI) [smets1995multidimensional], developed by Smets et al.; and the International Fitness Scale (IFIS) [Ortega2011], which measures changes in overall vigor and affect. These scales encompass a multitude dimensions of fatigue and serve as the foundation for traditional fatigue assessment. Nevertheless, both questionnaires and self-reports are highly influenced by an individual’s psychological state and external environment, resulting in subjective and inconsistent outcomes [gawron2016overview]. Moreover, most of these methods [brown2016fatigue, jap2009using, kar2010eeg] are post hoc assessments, making them unsuitable for real-time fatigue monitoring and preventive measures, both of which are essential in high-stakes environments, such as driving [app14073016] and aviation [article111, 10552405].
@@
Among various physiological signals, EEG has gained popularity in fatigue detection due to its ability to directly measure brain activity with high accuracy. However, traditional EEG systems are expensive, require an intrusive setup and specialized personnel for operation, making it challenging for real-world deployment. As a result, the majority of EEG-based fatigue studies remain confined to controlled laboratory settings [trejo2015eeg, cheng2011mental, 8982166, 10107629]. Although portable EEG devices have emerged for field applications, their data quality is not as good as that of laboratory-grade devices. Additionally, they are susceptible to external noise and interference, significantly impacting their stability and detection accuracy [larocco2020systemic, shashidhar2020real].
@@ II-B Data Synthesis Methods
Deep learning, on the other hand, has provided promising solutions to this challenge. Recurrent Neural Networks (RNNs), particularly Long Short-Term Memory (LSTM) networks, have proven effective in capturing long-term temporal dependencies in multivariate time-series data. Several studies have demonstrated their success in a variety of applications, such as GRU-D, which incorporates decay mechanisms and masking for clinical time-series data [che2016recurrent]; an LSTM-based imputation model for ECG signals that outperforms classical methods [verma2019accurate]; and BRITS, which enables bidirectional gradient-based imputation without strong distributional assumptions [cao2018brits]. In parallel, Convolutional Neural Networks (CNN) have shown great success in extracting local spatial and temporal features. Their strength has inspired hybrid models that combine CNNs with LSTMs to jointly leverage spatial and temporal dependencies for more accurate imputation and signal reconstruction [Eum2022Imputation]. Additionally, GANs have emerged as a leading approach for producing high-fidelity synthetic signals that are statistically indistinguishable from real ones. Examples can be found in [10239214][NIA2024110129].
@@ III-D Framework
$\mathcal{X}^{s-}=\mathcal{X}_{\text{T}}\setminus\mathcal{X}^{\cap},\qquad\mathcal{X}^{s+}:=\mathcal{X}_{s}\setminus\mathcal{X}^{\cap}$
@@
3:  Apply $g_{s}$ to impute the missing features $\mathcal{X}^{s+}$ in $\mathcal{D}_{T}$ based on $\mathcal{D}_{s}$, and update $\mathcal{D}_{T}$.
@@ IV-A Experimental Setup
For the experiments to be elaborated in a later subsection, we select three datasets containing diverse physiological signals, with some modalities overlapping across datasets and others being unique. Meanwhile, three widely used classification methods are chosen, utilizing both individual datasets and the enhanced target data for training. To minimize discrepancies and enhance the usability of these datasets, data pre-processing is required before prior to experimentation. To that end, we describe the selected datasets, implementation environment, and the data pre-processing steps in detail.
@@ IV-A2 Implementation Environment and Data Preprocessing
For all experiments described below, the learning parameters are initialized as listed in Table II.
@@ IV-B3 Impact of Cross-Domain Imputed Modalities on Classification Performance
To assess the robustness and generalizability of our approach across model architectures, we train four different classifiers in this experiment—MLP, LSTM, CNN1D, and a Transformer[9994611]—on each dataset variant. We assess performance using multiple metrics, including accuracy, binary F1, Area Under the ROC Curv(ROC AUC), Area Under the Precision–Recall Curve(PR AUC), macro F1, weighted F1, expected calibration error (ECE), and Brier score. This setup allows us to evaluate not only how each additional reconstructed modality—individually or in combination—contributes to downstream fatigue detection performance, but also the effectiveness of our framework across diverse classifier types.
@@ V Conclusion
Fatigue detection is crucial in high-stakes environments, where reduced alertness can have serious consequences. Despite advances in sensor technology, reliability and accessibility in real world environments remain major challenges, as many high-end sensors, such as EEG, are expensive, highly sensitive to environmental conditions, and require substantial operational setup outside controlled settings. To address this challenge, this paper proposes a proof-of-concept framework that enables knowledge transfer across differently instrumented sensor domains, aiming to enhance fatigue detection in real-world sensor-constrained scenarios. In particular, our approach transfers knowledge learned from high-fidelity laboratory sensors to field-deployable systems by using shared-channel overlap as a bridge, thereby maintaining detection accuracy when some modalities are unavailable. Our experiments provide empirical support for our claim and validate the framework’s effectiveness and practicality, showing that transferred knowledge can substantially enhance detection when high-end sensors are unavailable.
Although this paper focuses on the framework, our experiments reveal that incorporating more advanced algorithms can offer additional benefits. In particular, we observe that using TimeMixer and Transformer yields small but consistent gains over traditional configurations. We plan to explore this direction in future work by developing jointly optimized imputation and classification architectures to better capture nonlinear dependencies and enhance fatigue-detection performance. As the framework continues to evolve, we aim to further examine its real-time deployment aspects, focusing on computational efficiency, latency, and resource requirements to ensure suitability for field applications.